- mA 25 mV mA where r,- V/V A,=- =- =- V/v
Answer #1

1. 91-0-15mA LSK lc = α.一0.qq x o-75 ㎃ =0-146mA dl 25 mV 3320 where r, = stos 217 4-33.3 2. 230mA 2.0 =-20V/V